### v1

@@ 1 Introduction
Fault-tolerant quantum computation will be the basis of future large-scale quantum computers. To protect logical quantum information from environmental noise quantum error correction is essential. Among the array of promising quantum error-correcting codes the surface code [1, 2] stands out as a favored choice due to its high threshold, scalability, and planar connectivity. Recently, surface code memory experiments have been realized on various platforms [3, 4, 5, 6, 7], and the experimental investigation of fault-tolerant quantum gates has started [8].
@@ 2 The surface code and spacetime diagrams
In this paper we investigate the impact of errors on quantum operations between surface code patches – concepts that we briefly introduce in this Section.
@@ 2.1 A single patch of the surface code
$B_{f}$ $=\prod_{j\in\partial f}Z_{j};$ $A_{f}$ $=\prod_{j\in\partial f}X_{j}.$ (1)
@@
|$\ket{\overline{0}}\!$ $=\!\prod_{f\in\text{light}}\!\!\!\frac{1+A_{f}}{\sqrt{2}}\ket{0}^{\otimes n};$ |$\ket{\overline{1}}\!$ $=\!\prod_{f\in\text{light}}\!\!\!\frac{1+A_{f}}{\sqrt{2}}\ket{1}^{\otimes n}.$ (2)
Thus, $A_{f}\ket{\overline{\psi}}=B_{f}\ket{\overline{\psi}}=\ket{\overline{\psi}}$ for all logical states |$\ket{\overline{\psi}}$.
@@ 2.2 Errors, stabilizer measurements, check graphs
We consider two types of errors, against which the surface code protects the encoded qubit: data qubit errors and phenomenological readout errors. We restrict data qubit errors to probabilistic single-qubit Pauli $X$ and $Z$ (if these occur on the same qubit, they result in Pauli $Y$). Phenomenological readout errors are perfectly performed measurements with erroneously reported results. Both types of errors can be detected by repeated measurements of the stabilizer operators. We will refer to the errors as
@@
To handle these errors in the same manner we introduce virtual qubits representing stabilizer measurements. As shown in Fig. 3, virtual qubits (white dots) are placed between consecutive layers (time steps) of physical qubits (black dots). Spacelike errors are located on physical qubits and timelike errors are on virtual ones.
Errors are detected by check generators: these are assigned to the faces $f$ at time $t$, and are the products of the measurement outcomes of the corresponding stabilizer right before and after $t$. (The definition of check generators slightly differ at the boundaries which we will discuss later.) Thus, an odd number of data qubit/measurement errors result in a check generator value of $-1$, we call this a check violation. The locations of check violations contain the classical information that we gain from stabilizer measurements, this is called the syndrome.
The check graph contains the check generators as nodes and qubits (both physical and virtual) as edges. The check graph consists of two disconnected pieces: the $X$-check graph contains $X$-check generators (comparing measurement outcomes of $X$-stabilizers), and the $Z$-check graph contains $Z$-check generators. $X$-type errors can violate $Z$-check generators, while $Z$-type errors can violate $X$ ones. Notably, spacelike errors can be correlated due to Pauli $Y$ noise. However, timelike errors are not correlated, because readout errors on $Z$-stabilizers ($X$-type timelike errors) and readout errors on $X$-stabilizers ($Z$-type timelike errors) are independent.
To infer errors from the measurement data (the syndrome) a decoding algorithm has to be run on the check graph; we took the minimum-weight perfect matching (MWPM) [24, 25] approach. This produces a correction string for any given syndrome, which has both timelike and spacelike corrections. We used the efficient Blossom algorithm [26] to find the minimum-weight string, as implemented in the PyMatching software package [27, 28]. This solves the decoding on $X$- and $Z$-check graphs independently, with proper weighting of the edges to account for different data qubit and readout error rates. Our codes and the numerical data are available at [29]
@@ 2.3 Timelike and spacelike boundaries of check graphs
Boundaries are important features of check graphs. The $X$- and $Z$-boundaries, defined above for a single surface code patch, can be generalized for a check graphs – examples are shown in Fig. 3. Spacelike boundaries can be defined as surfaces containing the same boundary types in each stabilizer measurement round. Differently put, spacelike $X$- ($Z$-) boundaries are vertical surfaces, where only $X$- ($Z$-) check generators are present. Timelike boundaries can be defined in a similar manner, timelike $X$- ($Z$-) boundaries are horizontal surfaces, where only $X$- ($Z$-) check generators are present. We also define perfect timelike boundaries, where both $X$- and $Z$-check generators are present.
While timelike $X$- and $Z$-boundaries represent fault-tolerant initialization and measurement protocols, perfect timelike boundaries are just theoretical constructions. These are needed to describe the fault-tolerant logical operation itself, excluding errors occurring during initialization and during the final measurements.
Figure 3: Spacetime diagram segments and the corresponding check graphs representing the three different timelike boundaries. Black (white) dots: physical (virtual) qubits, blue (orange) volumes: $Z$- ($X$-) check generators. Top row: upper timelike $Z$-boundary, with a) spacetime diagram, b) $Z$-check graph, c) $X$-check graph. Here only $Z$-check generators are present at the top boundary. Middle row: upper timelike $X$-boundary, with d) spacetime diagram, e) $Z$-check graph, f) $X$-check graph. Here only $X$-check generators are present at the top boundary. Bottom row: upper perfect timelike boundary, with g) spacetime diagram, h) $Z$-check graph, j) $X$-check graph. These both contain check generators at the top boundary.
Initializing physical quantum bits in |$\ket{0}$ ($\ket{+}$) generates a timelike $Z$- ($X$-) boundary: in this case, the values of $Z$- ($X$-) stabilizers are predetermined before the first round of stabilizer measurements. Therefore, $Z$- ($X$-) check generators at such a boundary contain physical qubits from the "zeroth" time step and virtual qubits from the first stabilizer measurement round. In contrast, the measurement outcomes of $X$- ($Z$-) stabilizers are random, therefore there is a missing row of $X$- ($Z$-) check generators at this boundary. Similarly, measurement of physical qubits in the $Z$- ($X$-) basis generates timelike $Z$- ($X$-) boundaries. We illustrate the three different timelike boundaries in Fig. 3.
@@ 2.4 Spacetime diagrams: visualization of protocols
Logical protocols can often be visualized by so-called spacetime diagrams [30]. These do not include all details of the check graph representation, but contain the important features such as boundaries and twist defects, which separate different boundary types; examples are shown in Fig. 3. Spacetime diagrams are useful tools to identify different logical error strings and follow their propagation through the protocol.
@@ 4.1 Symmetry of the CNOT protocol and its consequences on the logical errors
Figure 8: Symmetry of the spacetime diagram of the CNOT. For equal $X$ and $Z$ error rates, the syndrome graph has a symmetry transformation which consists of the following steps: 1) time reversal, 2) mirror reflection, 3) swapping the $X$ and $Z$ labels (colors).
@@ 4.2 Numerical characterization of the logical error channel
Figure 10: Logical Pauli error rates in CNOT as a function of the physical noise parameter for patch size $d=7$, with (a) independent Pauli X and Z errors with equal error rates, and (b) depolarizing errors on the physical qubits. Numerically obtained error rates are shown with symbols with errorbars. Rates of symmetry partners have the same color; they are in most cases equal to good precision. Semi-theoretical expectations for these rates, calculated from $p_{XY}$, $p_{ZX}$, and $p_{YZ}$, as described in the main text, are shown with dashed lines, with shading indicating expected standard deviation.
@@
We found in all cases with independent noise that the logical error channel has the symmetry property expected based on the spacetime diagram, as per Fig. 9, to a good precision. A numerical example for patch size $d=5$ is also shown in Fig. 9. For a more systematic exploration, we 1) first extracted the string connection probability parameters $p_{1},p_{2},p_{3}$ by $p_{1}=\sqrt{p_{XY}}$, $p_{2}=\sqrt{p_{ZX}}$, $p_{3}=\sqrt{p_{YZ}}$, then 2) constructed the theoretical expectation for all Pauli error probabilities as per Fig. 9, then 3) compared the numerically obtained values with these predictions. We show this comparison for patch size $d=7$ in Fig. 10, and for other parameter sets in Appendix A.
@@ Appendix A Numerical results for the logical CNOT gate
Figure 12: Logical Pauli error rates in CNOT as a function of the physical noise parameter for various patch sizes $d$, with independent Pauli X and Z errors with equal error rates, or depolarizing errors on the physical qubits, as indicated on the plots. Numerically obtained error rates are shown with symbols with errorbars. Rates of symmetry partners have the same color; they are in most cases equal to good precision. Semi-theoretical expectations for these rates, calculated from $p_{XY}$, $p_{ZX}$, and $p_{YZ}$, as described in the main text, are shown with dashed lines, with shading indicating expected standard deviation. The other geometrical parameters of the procedure for all patch sizes are: $w=1$, $h_{1}=h_{3}=1$, $h_{2}=d$.
@@ Appendix B Correlations in the logical error channel under minimum weight perfect matching decoding
Eq.12 gives exactly zero for independent physical noise. Our numerical results for only spacelike errors ($d\times d$ case, with a single measurement round) and for spacelike and timelike errors ($d\times d\times d$ case, with $d$ measurement rounds) are presented in Fig. 13. Where it is visible that as the system increases, the correlation measure $M$ decreases. This effect is stronger for the $d\times d\times d$ case, because timelike $X$ and $Z$ errors are inherently independent, and the bulk is larger in this case. Therefore, we assume this effect only depends on the bulk size of the check graph rather than the actual shape. This effect explain our findings for the logical CNOT operation in the case of depolarizing noise.

### v3

@@ 1 Introduction
Fault-tolerant quantum computation will be the basis of future large-scale quantum computers. To protect logical quantum information from environmental noise quantum error correction is essential. Among the array of promising quantum error-correcting codes the surface code [1, 2] stands out as a favored choice due to its high threshold, scalability, and planar connectivity. Recently, surface code memory experiments have been realized on various platforms [3, 4, 5, 6, 7], and the experimental investigation of fault-tolerant quantum gates has started [8, 9, 10, 11].
@@ 2 The surface code and spacetime diagrams
In this paper we investigate the impact of errors on quantum operations between surface code patches – concepts that we briefly introduce in this Section. We also introduce the three-dimensional mathematical structure used for decoding, i.e., the check graph, and its simplified representation, the spacetime diagrams.
@@ 2.1 A single patch of the surface code
$B_{f}$ $=\prod_{j\in\partial f}Z_{j};$ $A_{f}$ $=\prod_{j\in\partial f}X_{j},$ (1)
@@
|$\ket{\overline{0}}\!$ $=\!\prod_{f\in\text{light}}\!\!\!\frac{1+A_{f}}{\sqrt{2}}\ket{0}^{\otimes n};$ |$\ket{\overline{1}}\!$ $=\!\prod_{f\in\text{light}}\!\!\!\frac{1+A_{f}}{\sqrt{2}}\ket{1}^{\otimes n},$ (2)
where the subscript "light" denotes faces with light shading, corresponding to $X$-stabilizers. Thus, $A_{f}\ket{\overline{\psi}}=B_{f}\ket{\overline{\psi}}=\ket{\overline{\psi}}$ for all logical states |$\ket{\overline{\psi}}$.
@@ 2.2 Errors, stabilizer measurements, check graphs
We consider two types of errors, against which the surface code protects the encoded qubit during a logical operation: data qubit errors and phenomenological readout errors. We restrict data qubit errors to probabilistic single-qubit Pauli $X$ and $Z$ (if these occur on the same qubit, they result in Pauli $Y$). Phenomenological readout errors are perfectly performed measurements with erroneously reported results. Both types of errors can be detected by repeated measurements of the stabilizer operators. We will refer to the errors as
@@
To detect errors, we need to repeatedly measure the $Z$- and $X$- stabilizers; this is represented by virtual bits. In Figs. 3, 4, the virtual bit obtained from a measurement at time $t$ is a white dot after the layer at time $t$ of physical qubits (black dots). Spacelike errors are located on physical qubits and timelike errors are on virtual bits (an erroneous stabilizer measurement is a bit-flip on a virtual bit).
The changes that errors cause in the measured Z- and X- stabilizer values are represented by check generators (also called detectors [27, 28]). These are generalizations of the concept of stabilizers : collections of measurement outcomes that should be compared to signal locations of errors. For intermediate time steps during logical operations, a check generator consists of consecutive stabilizer measurement outcomes, as in Fig.3, and its value is
$s_{f}^{t}=m_{f}^{t}+m_{f}^{t-1}\quad(\text{mod 2}),$ (3)
where $m_{f}^{t}$ is the outcome of the measurement of $A_{f}$ (or $B_{f}$) at time $t$. If data qubit measurements or initializations at time $t$ or $t-1$ are involved, the check generators slightly differs from Eq.(3), as we will discuss later.
Figure 3: $X$-type and $Z$-type check generators consist of the stabilizer measurements at time $t$ and $t-1$. The value is determined by the sum of the outcomes, as per Eq. (3), and can detect errors on the physical qubits and virtual bits.
All the information from the check generators obtained during a logical operation, performed using the surface code, is represented by a check graph. This graph has the check generators at the nodes and the possible errors (physical qubits and virtual bits) as edges. Each edge has a weight
$w_{e}=\ln\Big{(}\dfrac{1-p_{e}}{p_{e}}\Big{)},$ (4)
where $p_{e}$ is the probability of a spacelike/timelike error on the given data qubit/virtual bit.
The check graph consists of two disconnected pieces: the $X$-check graph contains $X$-check generators (comparing measurement outcomes of $X$-stabilizers), and the $Z$-check graph contains $Z$-check generators. $X$-type errors can violate $Z$-check generators, while $Z$-type errors can violate $X$ ones. Notably, correlations between $X$- and $Z$- spacelike errors can occur due to Pauli $Y$ noise. However, no such correlations occur on timelike errors, because readout errors on $Z$- and $X$- stabilizers ($X$- and $Z$- type timelike errors) are independent.
To infer errors from the measurement data in the check graph (the syndrome) a decoding algorithm has to be run on it; we took the minimum-weight perfect matching (MWPM) [29, 30] approach. This produces a correction string for any given syndrome, which has both timelike and spacelike corrections. We used the efficient Blossom algorithm [31] to find the minimum-weight string, as implemented in the PyMatching software package [32, 33]. This solves the decoding on $X$- and $Z$-check graphs independently, with proper weighting of the edges to account for different data qubit and readout error rates. Our codes and the numerical data are available at [34]
@@ 2.3 Timelike and spacelike boundaries of check graphs
We must pay special attention to (2-dimensional) boundaries of (3-dimensional) check graphs, both spacelike and timelike boundaries, both of which can be $X$-type or $Z$-type. These are generalizations of the (1-dimensional) $X$- and $Z$-boundaries of a single (2-dimensional) patch of surface code in Sect. 2.1. Examples are shown in Fig. 4.
Spacelike boundaries can be defined as surfaces containing the same boundary types in each stabilizer measurement round. Differently put, spacelike $X$- ($Z$-) boundaries are vertical surfaces, where only $X$- ($Z$-) check generators are present.
Figure 4: Spacetime diagram segments and the corresponding check graphs representing the three different timelike boundaries. Black (white) dots: physical (virtual) qubits (bits), blue (orange) volumes: $Z$- ($X$-) check generators. Top row: upper timelike $Z$-boundary, with a) spacetime diagram, b) $Z$-check graph, c) $X$-check graph. Here only $Z$-check generators are present at the top boundary. There is an extra layer of physical qubits on the $Z$-check graph due to the data qubit readout. Middle row: upper timelike $X$-boundary, with d) spacetime diagram, e) $Z$-check graph, f) $X$-check graph. Here only $X$-check generators are present at the top boundary. Bottom row: upper perfect timelike boundary, with g) spacetime diagram, h) $Z$-check graph, j) $X$-check graph. These both contain check generators at the top boundary.
Timelike lower boundaries, of X- or Z-type, occur in a check graph on areas where physical qubits are initialized, in |$\ket{0}$, or in $\ket{+}$, respectively. In this case, the values of $Z$- (respectively, $X$-) stabilizers are predetermined before the first round of stabilizer measurements. Therefore, $Z$- (respectively, $X$-) check generators at such a boundary contain only single stabilizer measurements, and their values are given as:
$s_{f}^{t_{0}}=m_{f}^{t_{0}},$ (5)
where the qubits were initialized directly before time $t_{0}$. In contrast, the measurement outcomes of $X$- (respectively, $Z$-) stabilizers are random, therefore there is a missing layer of $X$- (respectively, $Z$-) check generators at this boundary.
Timelike upper boundaries correspond to measurements of data qubits in some region in the $Z$- ( respetively, $X$-) basis at some time $t_{f}$. In that case the values of $Z$- (respectively, $X$-) stabilizers can be constructed from the data qubit measurements. Therefore, $Z$- (respectively, $X$-) check generators at such boundaries contain the stabilizer measurements from the $t_{f}-1$-th step, and also the corresponding data qubit measurements from the $t_{f}$-th step. The values of such check generators are given as:
$s_{f}^{t_{f}}=m_{f}^{t_{f}-1}+\sum_{j\in\partial f}m_{j}\quad(\text{mod 2}),$ (6)
where $m_{j}$-s are the data qubit measurement outcomes. Thus, similarly to lower timelike boundaries, these boundaries have an extra layer of $Z$-check generators with resepect to $X$-generators (respectively, extra layer of $X$- check generators).
Perfect timelike boundaries are timelike boundaries of check graphs with both check generator types present. This scenario cannot be constructed by any specific data qubit initialization/measurement pattern. However, this is a useful artificial construction for the investigation of fault-tolerant logical protocols.
We illustrate the three different timelike boundaries with the corresponding check graphs in Fig. 4.
@@ 2.4 Spacetime diagrams: visualization of protocols
For a simple overview of a realization of a logical operation using the surface code, one can represent only the boundaries of the check graph in a so-called spacetime diagram [35], as in Fig. 4. We will use this simplified representation often below, as spacetime diagrams are useful tools to identify different logical error strings and follow their propagation through the protocol.
@@ 4.1 Symmetry of the CNOT protocol and its consequences on the logical errors
Figure 9: Symmetry transformation of the spacetime diagram of the CNOT (a), in 3 steps. First, time reversal, resulting in (b); Then, mirror reflection, resulting in (c); finally, swapping the $X$ and $Z$ labels (colors), resulting in (d), identical to (a).
@@ 4.2 Numerical characterization of the logical error channel
Figure 11: Logical Pauli error rates in CNOT as a function of the physical noise parameter for patch size $d=7$, with (a) independent Pauli $X$ and $Z$ errors with equal error rates, and (b) depolarizing errors on the physical qubits. Numerically obtained error rates are shown with symbols with errorbars. Rates of symmetry partners have the same color; they are in most cases equal to good precision. Expectations for these rates, calculated from the fit parameters $p_{1},p_{2},p_{3}$, as described in the main text, are shown with dashed lines, with shading indicating expected uncertainty.
@@
We found in all cases with independent noise that the logical error channel has the symmetry property expected based on the spacetime diagram, as per Fig. 10, to a good precision. A numerical example for patch size $d=5$ is also shown in Fig. 10. For a more systematic exploration, we used a two-step procedure: 1) we first extracted the string connection probability parameters $p_{1}$, $p_{2}$, $p_{3}$ for each physical error probability by fitting a function in the form of Fig. 10: two-qubit Pauli errors as arguments, and $p_{1}$, $p_{2}$ and $p_{3}$ as fitting parameters (obtaining the logical channel with the structure of Fig. 10 closest to the numerical values); 2) we then compared the numerically obtained values of the frequency of two-qubit errors with the values obtained from the fitted model. We show this comparison for patch size $d=7$ in Fig. 11, and for other parameter sets in Appendix A.
@@ Appendix A Numerical results for the logical CNOT gate
Figure 13: Logical Pauli error rates in CNOT as a function of the physical noise parameter for various patch sizes $d$, with independent Pauli $X$ and $Z$ errors with equal error rates, or depolarizing errors on the physical qubits, as indicated on the plots. Numerically obtained error rates are shown with symbols with errorbars. Rates of symmetry partners have the same color; they are in most cases equal to good precision. Expectations for these rates, calculated from the fit parameters $p_{1},p_{2},p_{3}$, as described in the main text, are shown with dashed lines, with shading indicating expected uncertainty. The other geometrical parameters of the procedure for all patch sizes are: $w=1$, $h_{1}=h_{3}=1$, $h_{2}=d$.